OpenStudy (anonymous):

The equation r= 0.028t +20.8 can be used to predict the world record in the 200-m dash where R is the record in seconds and t is the number of years since 1920. In what year will the record be 18.0 sec?

OpenStudy (anonymous):

R= -0.028t + 20.8

OpenStudy (anonymous):

someone please help

OpenStudy (anonymous):

R= 18=-0.028t+20.8

OpenStudy (anonymous):

0.028t=20.8-18 t=(20.8-18)/0.028

OpenStudy (anonymous):

year=1920+t